A manufacturer of chocolate chips would like to know whether its bag filling machine works correctly at the 430 gram setting. It is believed that the machine is underfilling the bags. A 21 bag sample had a mean of 421 grams with a standard deviation of 15 15 . Assume the population is normally distributed. A level of significance of 0.1 will be used. Find the P-value of the test statistic. You may write the P-value as a range using interval notation, or as a decimal value rounded to four decimal places.
Answer #1

Answer:

\mu= 430, s=15, n=21, \bar x = 421, \alpha = 0.1

Ho: \mu \geq 430
Ha: \mu < 430 (claim)

formula for test statistics is

- s/n

+_421 - 430 15/V21

t= −2.75

Test Statistics = −2.75

Calculate P-Value for left tailed test with \alpha = 0.1

using t table we get P-Value as

P-value = 0.0062

since (P-Value= 0.0062) < ( \alpha = 0.1)

Reject the null hypothesis.